OpenStudy (kamille):

I need to solve this: arccos4x-2=0 @amistre64 could you please have a look?

OpenStudy (anonymous):

dude it is so simple place the 2 at the other hand

OpenStudy (anonymous):

and then solve for it

OpenStudy (amistre64):

is the -2 part of the arguement of the arccos?

OpenStudy (kamille):

no, but actually I made the problem incorrectly. It actually asks: When the equation has any solutions? (I need to find 'meanings' of a, if I understood correctly now) \[acos4x-2=0\]

OpenStudy (kamille):

I know that Ef of cosx must be [-1;1]

OpenStudy (kamille):

and only then it could have any solutions?

OpenStudy (amistre64):

not sure i understand the nature of the question yet. are you able to post it as a picture? to omit lost in translation errors?

OpenStudy (kamille):

\[-1\le a*4x-2<1\] ?

OpenStudy (kamille):

ah,but the question is in my own language, you can check, it is 4b http://i.imgur.com/ZQDo1uz.png

OpenStudy (kamille):

it says 'which meanings 'a' should have so equation has any solutions'

OpenStudy (amistre64):

google says its luthuainian

OpenStudy (kamille):

haha yes

OpenStudy (kamille):

sorry! anyways, thanks,I will try to do myself, I know the answer

OpenStudy (amistre64):

a cos(4x)-2 = 0 a cos(4x) = 2 a = 2 sec(4x)

OpenStudy (amistre64):

its just a simple algebra process

OpenStudy (kamille):

but the answer is \[a \in (-\infty;-2]U[2;+\infty)\]

OpenStudy (amistre64):

sec(4x) has some values of x to omit, namely some modified version of odd multiples of pi/2 attributed to the period adjustment

OpenStudy (amistre64):

sec(4x) ~~ cos(4x) = 0 when 4x = (2n+1)pi/2 x = (2n+1)pi/8

OpenStudy (amistre64):

so the range of 2 sec(4x) is the solution to a

OpenStudy (amistre64):

regardless of the period, the range is +- 2 to +- inf

OpenStudy (amistre64):